# Classical and quantum chaotic synchronization in coupled dissipative time crystals

### v4

@@ 2.1 Lindblad equation and classical mean field limit
If we consider the reduced variables $\hat{m}_{A/B}^{\alpha}=\hat{S}_{A/B}^{\alpha}/S$ we see that they commute in the limit $S\to\infty$. In this limit, therefore, one can neglect quantum correlations for these variables. So, rewriting Eq. (4) for the reduced variables, taking the limit $S\to\infty$ and neglecting quantum correlations, one obtains
@@ 3.2 Results
Figure 2: Largest Lyapunov exponent $\braket{\Lambda_{L}}$ (top row) and Pearson correlation coefficient $\overline{C_{P}}$ (bottom row) versus $\Gamma$ along the line $\Gamma+\Omega=1.5$ (left column), $2\Omega+\Gamma=2$ (center column), and $2\Omega+\Gamma=1.5$ (right column). These values are calculated as averages over $N_{\rm r}=100$ initial conditions generated on a spherical cap with parameter $a=0.1$.
@@
In Fig. 1(b), we consider three lines cutting the parameter space in Fig. 1(b) from the upper left to the bottom right part of the panel; These lines are $\Omega+\Gamma=1.5$, $2\Omega+\Gamma=2$, and $2\Omega+\Gamma=1.5$ and are marked in Fig. 1(b) with different symbols (blue circles, green triangles, and orange squares, respectively). Along these lines, we plot, using the same symbols, the LLE (Fig. 2, top row) and the corresponding average Pearson coefficient described in Sec. 3.1.4 (Fig. 2, bottom row) versus $\Gamma$. Both the LLE and the Pearson coefficient are evaluated over a set of random initial conditions with $a=0.1$. First of all, we notice in panels 2(c,d) that there is a regime of negative Lyapunov exponent corresponding to the trivial phase. Beyond that, we see that there is a threshold value $\Gamma\simeq 1.2$ in panel 2(a), $\Gamma\simeq 1.25$ in panel 2(b), $\Gamma\simeq 1.15$ in panel 2(c)] where the LLE moves from zero to positive, coinciding with a sudden increase of the Pearson coefficient. At this threshold, the Pearson coefficient shifts from strongly negative to strongly positive values. Below this threshold the system is in the CTC3 phase: Here, the Lyapunov exponent is vanishing, and the negative Pearson coefficient indicates that the two subsystem oscillate out of phase with respect to each other. Above the threshold the system becomes chaotic – as marked by the Lyapunov exponent becoming positive – but the two oscillators behave in a strictly correlated way, as witnessed by the positive value of the Pearson coefficient. Therefore, in this regime, the two subsystems display full chaotic synchronization.
@@
Figure 4: Time and initial-condition average values of $m_{j}^{z}(t)$, $\braket{\overline{m_{j}^{z}}}$ versus $\Gamma$, along the parameter-space line $\Omega+\Gamma=1.5$ (a, left column), $2\Omega+\Gamma=2$ (b, center column), and $2\Omega+\Gamma=1.5$ (c, right column). We average over a set of $N_{\rm r}=100$ initial conditions generated either for $a=0.1$ (top row) or $a=1$ (bottom row). The legend is shared. The background shadowing marks the Lyapunov exponent with the same color code as Fig. 1.
@@ 4 Quantum finite-size dynamics
Figure 5: (a,c) $\overline{S_{A}^{z}}$ and $\overline{S_{B}^{z}}$ versus $\Omega$ for different NN. (b,d) Time-averaged connected correlator $\overline{C_{AB}^{zz}}$. Average over $N_{\rm r}\geq 48$ trajectories. On the left column we consider the curves for the line $2\Omega+\Gamma=2$ and on the right one $2\Omega+\Gamma=1.5$. The vertical line marks the classical staggered-uniform crossover as found in Fig. 3(f).
@@
Figure 6: (a) Histogram of $(S_{A}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (b) Histogram of $(S_{B}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (c) Histogram of the entanglement entropy $\mathcal{E}^{r}(t)$ as it changes along time $t$ and across trajectories $r$. We fix the constraint $2\Omega+\Gamma=1.5$, $\delta t=10^{-4}$ and $S=20$.
We notice that the threshold for this crossover in the quantum case is in general not consistent with the mean field case, as we can see comparing Fig. 7(a,c) with Fig. 3(f). We see this fact comparing the curves with the vertical line in the plots that marks the classical crossover from the staggered-magnetization to the uniform-magnetization regime, as read from 3(f). While along the parameter-space line $2\Omega+\Gamma=2$ there is a reasonable correspondence (left column of Fig. 7), this is not true along a different parameter-space line ($2\Omega+\Gamma=1.5$ – left column of Fig. 7. Nevertheless, the quantum behavior is qualitatively similar to the classical one, because in both Fig. 4 and in Fig. 6 one sees a similar staggered-uniform crossover.
@@
for a time long enough that the steady state has been reached. This is the so-called nonequilibrium steady state (NESS) density matrix, $\hat{\rho}_{\rm NESS}$. After obtaining $\hat{\rho}_{\rm NESS}$ as an average over realizations at a time long enough that all finite-time transients have died out, we evaluate the average level spacing ratio $r_{S}$ of this matrix as described in Ref. [52]. Denoting $\rho_{\alpha}$ the eigenvalues of $\hat{\rho}_{\rm NESS}$ taken in increasing order, the average level spacing ratio is defined as
$r_{S}=\frac{1}{2S-1}\sum_{\alpha=1}^{2S-1}\frac{\min(\rho_{\alpha+1}-\rho_{\alpha},\rho_{\alpha+2}-\rho_{\alpha+1})}{\max(\rho_{\alpha+1}-\rho_{\alpha},\rho_{\alpha+2}-\rho_{\alpha+1})}\,.$ (23)
@@
Evaluating it, we always find $r_{s}\simeq 0.59$ (see Fig. 8), a value only slightly smaller than the random-matrix Gaussian unitary ensemble value $r_{\rm GUE}\simeq 0.5996$ resulting from a numerical fit in Ref. [53]. We can conclude that the steady-state Lindblad density matrix behaves as a random matrix, a behavior that corresponds to quantum chaos [52]. Beyond that, it displays a sudden crossover from staggered to uniform magnetization strictly reminiscent of the synchronization phenomenology observed in the mean field [see Fig. 7(a,c)]. Therefore, also in the quantum finite-$S$ case, we can identify this phenomenon as quantum chaotic synchronization.
@@
Figure 7: (a,c) Maximum points of the histograms of $S_{A}^{z}$ and $S_{B}^{z}$ along and across trajectories versus $\Gamma$. (b,d) Maximum point of the histogram of the entanglement entropy versus $\Gamma$. We fix the constraint $2\Omega+\Gamma=2$ in the left column and $2\Omega+\Gamma=1.5$ in the right column. We take also $\delta t=10^{-4}$. The vertical line marks the classical staggered-uniform crossover as found in Fig. 3(f).
@@
The fact that the quantum and the classical crossover do not coincide is related with the noncommutativity of the infinite-$S$ limit and the infinite time-limit. Indeed, the infinite-$S$ limit shows persistent (periodic or aperiodic) oscillations, while the finite-$S$ size case relaxes for $t\to\infty$ to a NESS density matrix for any arbitrary large but finite $S$ [39]. Here we are studying the properties of this NESS density matrix, which are in general different from the ones of the long-time behavior in the mean-field regime (see examples in Refs. [54, 43]). So it is remarkable that the crossover is present in both scenarios, despite the non-commutativity of the limits and their different long-time behaviors. Additionally, entanglement plays a significant role in the quantum case but is absent classically. We also note that the noncommutative nature of these limits helps explain why the largest Lyapunov exponent, shown in Fig. 2, exhibits intervals of zero (regular dynamics) and positive (chaotic dynamics), whereas the NESS density matrix displays signatures of quantum chaos throughout the same parameter range (see Fig. 8). This difference arises from the LLE assessing the long-time dynamics of the mean-field, infinite-$S$ limit, whereas the NESS characterizes the long-time limit of finite-size quantum dynamics – the infinite-$S$ limit being taken first in the former case.
Figure 8: Average level spacing ratio of the NESS density matrix versus $\Gamma$ for different values of $S$, compared with the average level spacing ratio for the Gaussian Unitary Ensemble as evaluated in Ref. [53]. The flatness of the curve and the agreement with the GUE value increase for increasing subsystem spin magnitude $S$. We fix the constraint $2\Omega+\Gamma=2$ and evaluate $\hat{\rho}_{\rm NESS}$ as $\hat{\rho}_{t}$ with $t=100$ in units of $1/\kappa$. In Eq. (22) we average over $N_{\rm r}=48$ quantum trajectories and take $\delta t=10^{-4}$.
@@ 5 Conclusions
We acknowledge P. Solanki for fruitful discussions. G. P. and A. R. acknowledge financial support from PNRR MUR Project No. PE0000023-NQSTI. We acknowledge computational resources from the CINECA award under the ISCRA initiative and from MUR, PON “Ricerca e Innovazione 2014-2020,” under Grant No. PIR01_\_00011 - (I.Bi.S.Co.). This work was further supported by the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733, by MUR Project No. CN_\_00000013-ICSC (P. L.), and by the QuantERA II Programme STAQS project, which received funding from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733 (P. L.). E. P. acknowledges financial support from the Erasmus+ AURORA ALLIANCE fellowship program. E. P. acknowledges support of the Czech Science Foundation (project 25-17472S). E. P. acknowledges project IGA PrF-2025-010.

### v5

# arXiv:2509.20922v5[quant-ph] 02 Mar 2026

@@ 2.1 Lindblad equation and classical mean field limit
If we consider the reduced variables $\hat{m}_{A/B}^{\alpha}=\hat{S}_{A/B}^{\alpha}/S$ we see that they commute in the limit $S\to\infty$. The key point is that the $\hat{m}_{j}^{\alpha}$ variables obey the commutation rule $\hat{m}_{j}^{\alpha},\hat{m}_{l}^{\beta}]=i\delta_{jl}\epsilon^{\alpha\beta\gamma}\hat{m}_{j}^{\gamma}/S$, so that in the limit $S\to\infty$ they commute and the dynamics is classical. In this limit, therefore, one can neglect quantum correlations for these variables. So, rewriting Eq. (4) for the reduced variables, taking the limit $S\to\infty$ and neglecting quantum correlations, one obtains
@@ 3.2 Results
Figure 2: Largest Lyapunov exponent $\braket{\Lambda_{L}}$ (top row) and Pearson correlation coefficient $\overline{C_{P}}$ (bottom row) versus $\Gamma$ along the line $\Gamma+\Omega=1.5$ (left column), $2\Omega+\Gamma=2\kappa$ (center column), and $2\Omega+\Gamma=1.5\kappa$ (right column). These values are calculated as averages over $N_{\rm r}=100$ initial conditions generated on a spherical cap with parameter $a=0.1$.
@@
In Fig. 1(b), we consider three lines cutting the parameter space in Fig. 1(b) from the upper left to the bottom right part of the panel; These lines are $\Omega+\Gamma=1.5\kappa$, $2\Omega+\Gamma=2\kappa$, and $2\Omega+\Gamma=1.5\kappa$ and are marked in Fig. 1(b) with different symbols (blue circles, green triangles, and orange squares, respectively). Along these lines, we plot, using the same symbols, the LLE (Fig. 2, top row) and the corresponding average Pearson coefficient described in Sec. 3.1.4 (Fig. 2, bottom row) versus $\Gamma$. Both the LLE and the Pearson coefficient are evaluated over a set of random initial conditions with $a=0.1$. First of all, we notice in panels 2(c,d) that there is a regime of negative Lyapunov exponent corresponding to the trivial phase. Beyond that, we see that there is a threshold value $\Gamma\simeq 1.2$ in panel 2(a), $\Gamma\simeq 1.25$ in panel 2(b), $\Gamma\simeq 1.15$ in panel 2(c)] where the LLE moves from zero to positive, coinciding with a sudden increase of the Pearson coefficient. At this threshold, the Pearson coefficient shifts from strongly negative to strongly positive values. Below this threshold the system is in the CTC3 phase: Here, the Lyapunov exponent is vanishing, and the negative Pearson coefficient indicates that the two subsystem oscillate out of phase with respect to each other. Above the threshold the system becomes chaotic – as marked by the Lyapunov exponent becoming positive – but the two oscillators behave in a strictly correlated way, as witnessed by the positive value of the Pearson coefficient. Therefore, in this regime, the two subsystems display full chaotic synchronization.
@@
Figure 4: Time and initial-condition average values of $m_{j}^{z}(t)$, $\braket{\overline{m_{j}^{z}}}$ versus $\Gamma$, along the parameter-space line $\Omega+\Gamma=1.5\kappa$ (a, left column), $2\Omega+\Gamma=2\kappa$ (b, center column), and $2\Omega+\Gamma=1.5\kappa$ (c, right column). We average over a set of $N_{\rm r}=100$ initial conditions generated either for $a=0.1$ (top row) or $a=1$ (bottom row). The legend is shared. The background shadowing marks the Lyapunov exponent with the same color code as Fig. 1.
@@ 4 Quantum finite-size dynamics
Figure 5: (a,c) $\overline{S_{A}^{z}}$ and $\overline{S_{B}^{z}}$ versus $\Omega$ for different NN. (b,d) Time-averaged connected correlator $\overline{C_{AB}^{zz}}$. Average over $N_{\rm r}\geq 48$ trajectories. On the left column we consider the curves for the line $2\Omega+\Gamma=2\kappa$ and on the right one $2\Omega+\Gamma=1.5\kappa$. The vertical line marks the classical staggered-uniform crossover as found in Fig. 3(f).
@@
Figure 6: (a) Histogram of $(S_{A}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (b) Histogram of $(S_{B}^{z})^{r}(t)$ as it changes along time $t$ and across trajectories $r$. (c) Histogram of the entanglement entropy $\mathcal{E}^{r}(t)$ as it changes along time $t$ and across trajectories $r$. We fix the constraint $2\Omega+\Gamma=1.5\kappa$, $\delta t=10^{-4}$ and $S=20$.
We notice that the threshold for this crossover in the quantum case is in general not consistent with the mean field case, as we can see comparing Fig. 7(a,c) with Fig. 3(f). We see this fact comparing the curves with the vertical line in the plots that marks the classical crossover from the staggered-magnetization to the uniform-magnetization regime, as read from 3(f). While along the parameter-space line $2\Omega+\Gamma=2\kappa$ there is a reasonable correspondence (left column of Fig. 7), this is not true along a different parameter-space line ($2\Omega+\Gamma=1.5\kappa$ – left column of Fig. 7. Nevertheless, the quantum behavior is qualitatively similar to the classical one, because in both Fig. 4 and in Fig. 6 one sees a similar staggered-uniform crossover.
@@
for a time long enough that the steady state has been reached. This is the so-called nonequilibrium steady state (NESS) density matrix, $\hat{\rho}_{\rm NESS}$. After obtaining $\hat{\rho}_{\rm NESS}$ as an average over realizations and over a time window $t_{\rm i},t_{\rm f}]$, choosing $t_{\rm i}$ such that all finite-time transients have died out, we evaluate the average level spacing ratio $r_{S}$ of this matrix as described in Ref. [43]. Denoting $\rho_{\alpha}$ the eigenvalues of $\hat{\rho}_{\rm NESS}$ taken in increasing order, the average level spacing ratio is defined as
$r_{S}=\frac{1}{2S-1}\sum_{\alpha=1}^{2S-1}\frac{\min(\rho_{\alpha+1}-\rho_{\alpha},\rho_{\alpha+2}-\rho_{\alpha+1})}{\max(\rho_{\alpha+1}-\rho_{\alpha},\rho_{\alpha+2}-\rho_{\alpha+1})}\,$, (23)
with the $\rho_{\alpha}$ taken in increasing order. Evaluating $r_{S}$ (see Fig. 8), we always find a value only slightly smaller than the random-matrix Gaussian unitary ensemble value $r_{\rm GUE}\simeq 0.5996$ resulting from a numerical fit in Ref. [1]. The reason behind the GUE statistics is that, due to the lack of time-inversion symmetry, there is in general no basis where it is fully real.
We emphasize that using the eigenvalues of $\hat{\rho}_{\rm NESS}$ – as we do – is equivalent to using the eigenvalues of $\hat{H}_{\rm eff}=-\ln\hat{\rho}_{\rm NESS}$ as done in [43]. The point is that the difference between nearby eigenvalues of $\hat{\rho}_{\rm NESS}$ is very small $\sim 1/(2S+1)\ll 1$ and one can perform a linear expansion
$|\rho_{\alpha+1}-\rho_{\alpha}|=|\operatorname{e}^{-H_{\alpha+1}}-\operatorname{e}^{-H_{\alpha}}|\simeq\operatorname{e}^{-H_{\alpha}}|H_{\alpha+1}-H_{\alpha}|\,.$ (24)
When evaluating the level spacing ratio the factors $\operatorname{e}^{-H_{\alpha}}$ simplify. It is the same principle why no unfolding is needed when evaluating the level spacing ratio. The approximation is very good already for $S=10$, and the two results agree within the errorbar, as we can see in Fig. 8.
@@
We can conclude that the steady-state Lindblad density matrix behaves as a random matrix, a behavior that corresponds to quantum chaos [43]. Beyond that, it displays a sudden crossover from staggered to uniform magnetization strictly reminiscent of the synchronization phenomenology observed in the mean field [see Fig. 7(a,c)]. Therefore, also in the quantum finite-$S$ case, we can identify this phenomenon as quantum chaotic synchronization.
@@
Figure 7: (a,c) Maximum points of the histograms of $S_{A}^{z}$ and $S_{B}^{z}$ along and across trajectories versus $\Gamma$. (b,d) Maximum point of the histogram of the entanglement entropy versus $\Gamma$. We fix the constraint $2\Omega+\Gamma=2\kappa$ in the left column and $2\Omega+\Gamma=1.5\kappa$ in the right column. We take also $\delta t=10^{-4}$.
The fact that the quantum and the classical crossover do not coincide is related with the noncommutativity of the infinite-$S$ limit and the infinite time-limit. Indeed, the infinite-$S$ limit shows persistent (periodic or aperiodic) oscillations, while the finite-$S$ size case relaxes for $t\to\infty$ to a NESS density matrix for any arbitrary large but finite $S$ [49]. Indeed, our system has a well-defined semiclassical limit obtained by taking $S\to\infty$. In this limit, the mean-field equations exhibit a rich phenomenology at long times: convergence to stationary values (stationary phase), time-periodic oscillations (time-crystal phase), or irregular dynamics (chaotic phase).
Non trivial long-time dynamics can emerge only when the thermodynamic limit is taken first. In this scenario ($\lim_{t\to\infty}\lim_{S\to\infty}$), the mean-field system is described by the classical nonlinear equations Eq. (5) that can sustain persistent oscillations or chaos. By contrast, for any finite $S$, quantum fluctuations and the finite-dimensional Hilbert space cause all dynamics to eventually converge to a steady state, after a transient behavior. Taking the long-time limit first ($\lim_{S\to\infty}\lim_{t\to\infty}$) thus washes out the interesting dynamical features before one moves to the thermodynamic limit.
This subtle issue regarding the non-commutativity of limits has been discussed in recent literature on dissipative quantum systems. One example is Ref. [9], which analyzes dissipative quantum models that like ours are described by classical equations in the thermodynamic limit, and explicitly demonstrates how exchanging the order of these limits affects the phase diagram. In another work, quantum chaos in the driven-dissipative Bose-Hubbard dimer does not match the long-time regular behavior found in the classical limit [PhysRevResearch.7.013276]. Also this lack of quantum-classical correspondence can be interpreted as a noncommutativity of the thermodynamic and the long-time limits. Therefore, quantitative agreement remains inherently limited by the fundamental non-commutativity of the two limits.
@@
Here we are studying the properties of the finite-$S$ NESS density matrix, which – consistently with the discussion above – are in general different from the ones of the long-time behavior in the $S\to\infty$ mean-field regime [9, PhysRevResearch.7.013276]. So it is remarkable that the crossover is present in both scenarios, despite the non-commutativity of the limits and their different long-time behaviors. Additionally, entanglement plays a significant role in the quantum case but is classically absent.
We also note that the noncommutative nature of these limits helps explain why the largest Lyapunov exponent, shown in Fig. 2, exhibits intervals of zero (regular dynamics) and positive (chaotic dynamics), whereas the NESS density matrix displays signatures of quantum chaos throughout the same parameter range (see Fig. 8). This difference arises from the LLE assessing the long-time dynamics of the mean-field, infinite-$S$ limit, whereas the NESS characterizes the long-time limit of finite-size quantum dynamics – the infinite-$S$ limit being taken first in the former case. We emphasize that a similar behavior has been observed in a different model in [PhysRevResearch.7.013276].
Figure 8: The average level spacing ratio evaluated – as explained in the text – for the eigenvalues of $\hat{\rho}_{\rm NESS}$ and $\hat{H}_{\rm eff}$, and plotted versus $\Gamma$ along the line $2\Gamma+\Omega=2\kappa$, taking different values of $S$. The $\hat{\rho}_{\rm NESS}$ is evaluated using Eq. (22), averaging over $N_{\rm r}=48$ realizations and over a time window from $t_{\rm i}=80$ to $t_{\rm f}=100$, and then the average level spacing ratio $r_{S}$ is evaluated. $r_{S}$ is further averaged over $N_{\rm s}=10$ different sets of disorder realizations, and the errorbar is evaluated as the root mean square deviation over sets divided by $\sqrt{N}_{\rm s}$.
@@ 5 Conclusions
We acknowledge P. Solanki for fruitful discussions. G. P. and A. R. acknowledge financial support from PNRR MUR Project No. PE0000023-NQSTI. We acknowledge computational resources from the CINECA award under the ISCRA initiative and from MUR, PON “Ricerca e Innovazione 2014-2020,” under Grant No. PIR01_\_00011 - (I.Bi.S.Co.). This work was supported by PNRR MUR Project No. PE0000023-NQSTI, by the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733, by MUR Project No. CN_\_00000013-ICSC (P. L.), and by the QuantERA II Programme STAQS project, which received funding from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733 (P. L.). E. P. acknowledges financial support from the Erasmus+ AURORA ALLIANCE fellowship program. E. P. acknowledges support of the Czech Science Foundation (project 25-17472S). E. P. acknowledges project IGA PrF-2025-010. E. P. acknowledges use of the computational cluster of the Department of Optics, Palacký University.